OpenStudy (anonymous):

Which is not a solution to the equation ? x-2y=10 A. (–6, –8) B. (–3, 4) C. (8, –1) D. (14, 2)

OpenStudy (texaschic101):

we got to sub our answer choices in and see which one makes the equation true x - 2y = 10 (-6,-8) -6 - 2(-8) = 10 -6 + 16 = 10 10 = 10 (correct) I lucked out...it happened to be the first one...I didn't have to go through all of them....lol

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (texaschic101):

STOP...wait.....I messed up

OpenStudy (anonymous):

OK

OpenStudy (texaschic101):

Its asking for one that is not a solution....I did one that was a solution.. x - 2y = 10 (-3,4) -3 - 2(4) = 10 -3 - 8 = 10 - 11 = 10 Thats it....B.....sorry bout that

OpenStudy (anonymous):

iTS OK AND THANKS

OpenStudy (texaschic101):

:)